Suppose two sides of a square are given as x + 3 and 12 - x. What is its area?
Answer by jim_thompson5910(35256)

If this is a square, then it has equal sides. So x+3 = 12-x

Solve for x

x+3 = 12-x

x+x = 12-3

2x = 9

x = 9/2

x = 4.5

So each side is 4.5 + 3 = 7.5 units.


So the area is 7.5^2 = 56.25 square units.
------------------------------------------------------------------
If this is really a rectangle, then the area is simply

A = (x+3)(12-x)

A = 12x - x^2 + 36 - 3x

A = -x^2 + 9x + 36